# CBGBench: Fill in the Blank of Protein-Molecule Complex Binding Graph

### v2

@@ Abstract
Structure-based drug design (SBDD) aims to generate potential drugs that can bind to a target protein and is greatly expedited by the aid of AI techniques in generative models. However, a lack of systematic understanding persists due to the diverse settings, complex implementation, difficult reproducibility, and task singularity. Firstly, the absence of standardization can lead to unfair comparisons and inconclusive insights. To address this dilemma, we propose CBGBench, a comprehensive benchmark for SBDD, that unifies the task as a generative heterogeneous graph completion, analogous to fill-in-the-blank of the 3D complex binding graph. By categorizing existing methods based on their attributes, CBGBench facilitates a modular and extensible framework that implements various cutting-edge methods. Secondly, a single task on de novo molecule generation can hardly reflect their capabilities. To broaden the scope, we have adapted these models to a range of tasks essential in drug design, which are considered sub-tasks within the graph fill-in-the-blank tasks. These tasks include the generative designation of de novo molecules, linkers, fragments, scaffolds, and sidechains, all conditioned on the structures of protein pockets. Our evaluations are conducted with fairness, encompassing comprehensive perspectives on interaction, chemical properties, geometry authenticity, and substructure validity. We further provide the pre-trained versions of the state-of-the-art models and deep insights with analysis from empirical studies. The codebase for CBGBench is publicly accessible at https://github.com/Edapinenut/CBGBench.
@@ 1 Introduction
The rapid and remarkable progress in Graph Neural Networks (GNNs) and generative models have advanced the bio-molecule design in these years [1, 2, 3]. In structure-based drug design, AI-aided methods aim to learn the chemical space of the molecules that can bind to certain proteins as targets, which decreases the general chemical space of molecule compounds ($\sim 10^{60}$) to a more compact search space and enables the model to explore the potential binding drugs. Recent success in generative models such as diffusion models [4, 5, 6] further enhanced these methods to fully exploit the targeted chemical space, and AI-aided SBDD has been propelled into another prominence.
@@
Despite the significance of the SBDD and the development of various approaches, there remains a lack of a comprehensive benchmark for this field covering various practical application scenarios. On the one hand, although different methods are proposed for the task, the experimental setup is not unified, and the evaluation protocol also differs. For example, GraphBP [7] use a different training and test split of Crossdocked2020 [8] from concurrent works like Pocket2Mol [9] and 3DSBDD [10]; And DiffBP [11] and GraphBP employ Gnina [12] instead of AutoDock Vina [13] as the evaluator for obtaining docking score as binding affinity. Besides, the complex implementation of models makes the modules coupled, so it is hard to figure out which proposed module contributes to the improvements in performance. For example, though TargetDiff [14] and DiffBP are both diffusion-based methods, the first one uses EGNN [15] to predict the ground-truth position of each atom in the molecule while the later one uses GVP [16] to remove the added noise in a score-based way. To have a systematic understanding of the designation for diffusion models in molecule generation, the network architectures should be fixed to keep the expressivity equal. On the other hand, the task of de novo generation of molecules is a branch of SBDD. There are also other important tasks, such as the optimization of molecular side chains in lead optimization, or linker design to combine functional fragments into a single, connected molecule [17, 18]. It is highly meaningful to explore whether these methods can be successfully transferred to applications in drug optimization.
@@
In this way, here we propose CBGBench as a benchmark for SBDD, with the latest state-of-the-art methods included and integrated into a single codebase (See Appendix. B for details.). Firstly, we unify the generation of the binding molecule as a heterogeneous graph completion task, i.e. fill-in-the-blank of the 3D Complex Binding Graph. In this way, the systematic categorization of existing methods is based on three dichotomies: (i) voxelized v.s. continuous position generation, (ii) one-shot v.s. auto-regressive generation, and (iii)domain-knowledge-based v.s. full-data-driven generation. As a result, the methods can be easily modularized and extensible in a unified framework. Secondly, CBGBench introduces a unified protocol with a comprehensive evaluation. In chemical property, the classical ones including QED, SA, LogP and LPSK are considered; For interaction, besides Vina enery, we extend the evaluation to interaction types between protein and molecules; For geometry, we considered static ones including bond length and bond angles, and clashes resulted from protein-molecule cross-clash between atoms in the generated molecules; In substructure analysis, we consider single atoms, rings, and functional groups as pharmacophores. Moreover, thanks to our reformulation of the problem and extensible modular implementation, we can easily extend these methods to the other four tasks in lead optimization, including the designation of (i) linkers, (ii) fragments, (iii) side chains and (iv) scaffolds. Comprehensive evaluation is also conducted for these tasks, to explore the potential application value of the existing methods in lead optimization.
@@
  • CNN-based Methods by using density maps as molecule representations remain highly competitive in target-aware molecular generation.
@@
  • Most evaluated methods can be well generalized to lead optimization. Empirical studies show that scaffold hopping is the most challenging task among them, while linker design is relatively the easiest.
  • The conclusions drawn from the experimental results and with the evaluation protocol of CBGBench, are mainly consistent with those obtained on real-world targets, demonstrating generalizability of our benchmark.
@@ 2.1 Problem Statement
For a binding system composed of a protein-molecule pair as $(\mathcal{P},\mathcal{M})$, in which $\mathcal{P}$ contains $N_{\mathrm{rec}}$ atoms of proteins and $\mathcal{M}$ contains $N_{\mathrm{lig}}$ atoms of molecules, we represent the index set of the protein’s atom as $\mathcal{I}_{\mathrm{rec}}$ and the molecule’s atoms as $\mathcal{I}_{\mathrm{lig}}$, with $|\mathcal{I}_{\mathrm{rec}}|=N_{\mathrm{rec}}$ and $|\mathcal{I}_{\mathrm{lig}}|=N_{\mathrm{lig}}$. The binding graph can be regarded as a heterogenous graph. One subgraph is protein structures, as $\mathcal{P}=(\bm{V}_{\mathrm{rec}},\bm{E}_{\mathrm{rec}})$, where $\bm{V}_{\mathrm{rec}}=\{(a_{i},\bm{\mathrm{x}}_{i},s_{i})\}_{i\in\mathcal{I}_{\mathrm{rec}}}$ is the node set, and $\bm{E}_{\mathrm{pa}}=\{(i,j,e_{i,j})\}_{i,j\in\mathcal{I}_{\mathrm{rec}}}$ is the edge set. Here, $a_{i}$ is the atom types with $a_{i}=1,\ldots,M$, $\bm{\mathrm{x}}_{i}\in\mathbb{R}^{3}$ is the corresponding 3D position, and $s_{i}=1,\ldots,20$ is the amino acid types that $i$-th atom belongs to; elments in edge set means there exists a bond between $i$ and $j$, with edge type $e_{i,j}$. The other graph is the molecule structures, written as $\mathcal{M}=(\bm{V}_{\mathrm{lig}},\bm{E}_{\mathrm{lig}})$, where $\bm{V}_{\mathrm{lig}}=\{(a_{i},\bm{\mathrm{x}}_{i})\}_{i\in\mathcal{I}_{\mathrm{lig}}}$ and $\bm{E}_{\mathrm{lig}}$ is the edge set with the same form as $\bm{E}_{\mathrm{rec}}$. Besides, there are virtual edges that exist between protein and molecule, which make up cross edge set $\bm{E}_{\mathrm{crs}}$. Denote the probabilistic model for binding graph by $p(\mathcal{M},\mathcal{P})$. For the de novo molecule generation, the model aims to learn the probability of $p(\mathcal{M}|\mathcal{P})$, as to fill the blank of the protein pocket with a molecule, by using a generative model $p(\cdot)$.
By this means, for the other missing blanks, we write the generation node set as $\mathcal{G}=(\bm{V}_{\mathrm{gen}},\bm{E}_{\mathrm{gen}})$, and the molecule context set as $\mathcal{C}=(\bm{V}_{\mathrm{ctx}},\bm{E}_{\mathrm{ctx}})$, in which $\mathcal{I}_{\mathrm{ctx}}=\mathcal{I}_{{\mathrm{lig}}}\setminus\mathcal{I}_{{\mathrm{gen}}}$. Therefore, the generative model for filling the blank of parts of molecules aims to learn the probability of $p(\mathcal{G}|\mathcal{C},\mathcal{P})$. The task of fill-in-the-partial-graph is of great significance in SBDD, especially in lead optimization, and we establish other four sub-tasks of it besides de novo design, with accessible datasets in Sec. 3.
@@ 2.2 Related Work and Taxonomy
The SBDD methods are initially combined with deep neural networks on voxel grids, such as LiGAN [19] generating atom voxelized density maps using VAE [20], and 3DSBDD [10] predicts whether the position on grids is occupied with which type of atoms using auto-regressive models. Then, the development of Equivariant Graph Neural Networks (EGNNs) boosts the SBDD methods to directly generate the continuous 3D positions, such as Pocket2Mol [9] generating the molecules’ atom types, positions, and the connected bonds and GraphBP [7] employing normalizing flows to generate these attributes, both in an auto-regressive way. The diffusion models [4] further propel the AI-aided SBDD methods, such as TargetDiff [14], DiffBP [11] and DiffSBDD [21], generating full atoms’ positions and element types. Recently, Domain knowledge has been used to constrain or guide the generation of binding drugs. For example, D3FG [22] and FLAG [23] use prior knowledge of fragment motifs to model the coarse structures, and DecompDiff [24] harnesses virtual points searching for arm- and scaffold-clustering as prior knowledge with validity guidance for sampling.
In this way, we can categorize these methods with the three standards:
  • Whether the generation process is auto-regressive or one-shot.
To better articulate, we classify them according to Table. 1, and Figure. 1 gives a simple demonstration of the criteria for taxonomy.
Table 1: Categorization of included methods. Others except for LiGAN and 3DSBDD model atom positions as continuous variables; Diffusion-based methods maintain a constant atom number in a generation, as one-shot methods; FLAG and D3FG use fragment motifs, and DecompDiff uses arm&scaffolding priors with validity guidance as domain knowledge.
@@ 3 Task and Dataset
Besides the de novo molecule generation, the modularized methods can be easily extended to four subtasks which are branches of our defined fill-in-the-blank of the complex binding graph, including the target-aware linker, fragment, side chain, and scaffold design [25], leading to the generative targets $\mathcal{G}$ for the probabilistic model $p(\mathcal{G}|\mathcal{C},\mathcal{P})$ to be linkers, fragments, side chains, and scaffolds, and the molecular context $\mathcal{C}$ as the rest. We demonstrate the significance of the four tasks in lead optimization and how the corresponding datasets are established below. The datasets are all established based on previous splits of Crossdocked2020 [8] for fair comparison.
Linker Design. Linker design is a critical strategy in fragment-based drug discovery [26]. It focuses on creating linkers that connect two lead fragments and obtaining a complete molecule with enhanced affinity. Effective linker design can significantly influence the efficacy and pharmacokinetics of the drug, by ensuring that the molecule maintains the desired orientation and conformation when bound to its target [27]. We define linkers following specific rules: (i) A linker should act as the intermediate structure connecting two fragments. (ii) A linker should contain at least two atoms on the shortest path between two fragments. (iii) Each connecting fragment must consist of more than five atoms.
Fragment Growing. Fragment growing also plays an important role in fragment-based drug design, focusing on expanding a fragment on the lead compound to better fill the binding pocket [28]. It also relates to adjusting pharmacological properties, such as enhancing solubility or reducing toxicity [29]. In this approach, we decompose the entire ligand into two fragments and select the smaller one for expansion based on specific criteria: (i) Each fragment must contain more than five atoms. (ii) The smaller fragment should be larger than half the size of the larger one.
Side Chain Decoration. Side chain decoration overlaps with fragment growing but differs in that it permits modifications at multiple sites on the lead compound, whereas fragment growing typically modifies a single site. The taxonomy of arms-scaffold in some previous work [30] is similar to side-chain-scaffold, while we focus on a chemist-intuitive approach, Bemis-Murko decomposition [31], which treats all terminal non-cyclic structures as side chains.
Scaffold Hopping. Scaffold hopping is introduced by Schneider [32], as a strategy in medicinal chemistry aiming to replace the core structure of molecules to explore more diverse chemical structures or improve specific properties. While various empirical definitions exist [33], for consistency, the Bemis-Murko decomposition is utilized to define the scaffold structure.

### v3

@@ Abstract
Structure-based drug design (SBDD) aims to generate potential drugs that can bind to a target protein and is greatly expedited by the aid of AI techniques in generative models. However, a lack of systematic understanding persists due to the diverse settings, complex implementation, difficult reproducibility, and task singularity. Firstly, the absence of standardization can lead to unfair comparisons and inconclusive insights. To address this dilemma, we propose CBGBench, a comprehensive benchmark for SBDD, that unifies the task as a generative graph completion, analogous to fill-in-the-blank of the 3D complex binding graph. By categorizing existing methods based on their attributes, CBGBench facilitates a modular and extensible framework that implements various cutting-edge methods. Secondly, a single de novo molecule generation task can hardly reflect their capabilities. To broaden the scope, we adapt these models to a range of tasks essential in drug design, considered sub-tasks within the graph fill-in-the-blank tasks. These tasks include the generative designation of de novo molecules, linkers, fragments, scaffolds, and sidechains, all conditioned on the structures of protein pockets. Our evaluations are conducted with fairness, encompassing comprehensive perspectives on interaction, chemical properties, geometry authenticity, and substructure validity. We further provide deep insights with analysis from empirical studies. Our results indicate that there is potential for further improvements on many tasks, with optimization in network architectures, and effective incorporation of chemical prior knowledge. Finally, to lower the barrier to entry and facilitate further developments in the field, we also provide a single codebase (https://github.com/EDAPINENUT/CBGBench) that unifies the discussed state-of-the-art models, data pre-processing, training, sampling, and evaluation.
@@ 1 Introduction
The rapid and remarkable progress in Graph Neural Networks (GNNs) and generative models have advanced the bio-molecule design in these years (Jumper et al., 2021; Watson et al., 2023; Krishna et al., 2023). In structure-based drug design, AI-aided methods aim to learn the chemical space of the molecules that can bind to certain proteins as targets, which decreases the general chemical space of molecule compounds ($\sim 10^{60}$) to a more compact search space and enables the model to explore the potential binding drugs. Recent success in generative models such as diffusion models (Ho et al., 2020; Lipman et al., 2022; Song et al., 2020) further enhanced these methods to fully exploit the targeted chemical space, and AI-aided SBDD has been propelled into another prominence.
@@
Despite the significance of the SBDD and the development of various approaches, there remains a lack of a comprehensive benchmark for this field covering various practical application scenarios. On the one hand, although different methods are proposed for the task, the experimental setup is not unified, and the evaluation protocol also differs. For example, GraphBP (Liu et al., 2022) use a different training and test split of Crossdocked2020 (Francoeur et al., 2020) from concurrent works like Pocket2Mol (Peng et al., 2022) and 3DSBDD (Luo et al., 2022); And DiffBP (Lin et al., 2022) and GraphBP employ Gnina (McNutt et al., 2021) instead of AutoDock Vina (Trott & Olson, 2009) as the evaluator for obtaining docking score as binding affinity. Besides, the complex implementation of models makes the modules coupled, so it is hard to figure out which proposed module contributes to the improvements in performance. For example, though TargetDiff (Guan et al., 2023b) and DiffBP are both diffusion-based methods, the first one uses EGNN (Satorras et al., 2022) to predict the ground-truth position of each atom in the molecule while the later one uses GVP (Jing et al., 2020) to remove the added noise in a score-based way. To have a systematic understanding of the designation for diffusion models in molecule generation, the network architectures should be fixed to keep the expressivity equal. On the other hand, the task of de novo generation of molecules is a branch of SBDD. There are also other important tasks, such as the design of molecular side chains in lead optimization, or linker design to combine functional fragments into a single, connected molecule (Igashov et al., 2022; Guan et al., 2023a). It is highly meaningful to explore whether these methods can be successfully transferred to applications in drug optimization.
@@
In this way, here we propose CBGBench as a benchmark for SBDD, with the latest state-of-the-art methods included and integrated into a single codebase (See Appendix. C for details). Firstly, we unify the generation of the binding molecule as a graph completion task, i.e. fill-in-the-blank of the 3D Complex Binding Graph. In this way, the systematic categorization of existing methods is based on three dichotomies: (i) voxelized v.s. continuous position generation, (ii) one-shot v.s. auto-regressive generation, and (iii) domain-knowledge-based v.s. full-data-driven generation. As a result, the methods can be easily modularized and extensible in a unified framework. Secondly, CBGBench introduces a unified protocol with a comprehensive evaluation. In chemical property, the classical ones including QED, SA, LogP and LPSK are considered; For interaction, besides Vina enery, we extend the evaluation to interaction types between protein and molecules; For geometry, we considered static ones including bond length and bond angles, and clashes resulted from protein-molecule cross-clash between atoms in the generated molecules; In substructure analysis, we consider single atoms, rings, and functional groups as pharmacophores. Moreover, thanks to our reformulation of the problem and extensible modular implementation, we can easily extend these methods to the other four tasks in lead optimization, including the designation of (i) linkers, (ii) fragments, (iii) side chains and (iv) scaffolds. Comprehensive evaluation is also conducted for these tasks, to explore the potential application value of the existing methods in lead optimization.
@@
  • CNN-based methods by modeling voxelized density maps as molecule representations remain highly competitive in target-aware molecule generation, and the diffusion-based ones achieve the state-of-the-art.
@@
  • Most evaluated methods can be well generalized to lead optimization. Empirical studies show that scaffold hopping is the most challenging task among them, while linker design is relatively the easiest. However, there is still a large space for improvements in these tasks.
  • The conclusions drawn from the experimental results and with the evaluation protocol of CBGBench, are mainly consistent with those obtained on real-world targets, demonstrating generalizability and effectiveness of our benchmark.
@@ 2.1 Problem Statement
For a binding system composed of a protein-molecule pair as $(\mathcal{P},\mathcal{M})$, in which $\mathcal{P}$ contains $N_{\mathrm{rec}}$ atoms of proteins and $\mathcal{M}$ contains $N_{\mathrm{lig}}$ atoms of molecules, we represent the index set of the protein’s atom as $\mathcal{I}_{\mathrm{rec}}$ and the molecule’s atoms as $\mathcal{I}_{\mathrm{lig}}$, with $|\mathcal{I}_{\mathrm{rec}}|=N_{\mathrm{rec}}$ and $|\mathcal{I}_{\mathrm{lig}}|=N_{\mathrm{lig}}$. The binding graph can be regarded as a heterogenous graph. One subgraph is protein structures, as $\mathcal{P}=(\bm{V}_{\mathrm{rec}},\bm{E}_{\mathrm{rec}})$, where $\bm{V}_{\mathrm{rec}}=\{(a_{i},\bm{\mathrm{x}}_{i},s_{i})\}_{i\in\mathcal{I}_{\mathrm{rec}}}$ is the node set, and $\bm{E}_{\mathrm{rec}}=\{(i,j,e_{i,j})\}_{i,j\in\mathcal{I}_{\mathrm{rec}}}$ is the edge set. Here, $a_{i}$ is the atom types with $a_{i}=1,\ldots,M$, $\bm{\mathrm{x}}_{i}\in\mathbb{R}^{3}$ is the corresponding 3D position, and $s_{i}=1,\ldots,20$ is the amino acid types that $i$-th atom belongs to; elements in edge set means there exists
Figure 1: Filling the blank in the complex binding site.
a bond between $i$ and $j$, with edge type $e_{i,j}$. The other graph is the molecule structures, written as $\mathcal{M}=(\bm{V}_{\mathrm{lig}},\bm{E}_{\mathrm{lig}})$, where $\bm{V}_{\mathrm{lig}}=\{(a_{i},\bm{\mathrm{x}}_{i})\}_{i\in\mathcal{I}_{\mathrm{lig}}}$ and $\bm{E}_{\mathrm{lig}}$ is the edge set with the same form as $\bm{E}_{\mathrm{rec}}$. Besides, there are virtual edges that exist between protein and molecule, which make up cross-edge set $\bm{E}_{\mathrm{crs}}$. Denote the probabilistic model for binding graph by $p(\mathcal{M},\mathcal{P})$. For the de novo molecule generation, the model aims to learn the probability of $p(\mathcal{M}|\mathcal{P})$, as to fill the blank of the protein pocket with atoms in a molecule, by using a generative model $p(\cdot)$.
A comparable analogy can be made that if we also divide the atoms in a molecule into known and to-be-generated parts, we can extend the aforementioned ‘filling-in-the-blank’ task as shown in Figure. 1. In specific, for the missing blanks, we write the node set for generation as $\mathcal{G}=(\bm{V}_{\mathrm{gen}},\bm{E}_{\mathrm{gen}})$, and the known atoms in the molecule as elements in the context set $\mathcal{C}=(\bm{V}_{\mathrm{ctx}},\bm{E}_{\mathrm{ctx}})$, in which $\mathcal{I}_{\mathrm{ctx}}=\mathcal{I}_{{\mathrm{lig}}}\setminus\mathcal{I}_{{\mathrm{gen}}}$. Therefore, the generative model for filling the blanks of the missing parts of molecules aims to learn the probability of $p(\mathcal{G}|\mathcal{C},\mathcal{P})$. The newly-defined task of fill-in-the-partial-graph is of great significance in SBDD, especially in lead optimization, and we establish other four sub-tasks of it besides de novo design, with accessible datasets in Sec. 3.
@@ 2.2 Related Work and Taxonomy
The SBDD methods are initially combined with deep neural networks on voxel grids, such as LiGAN (Masuda et al., 2020) generating atom voxelized density maps using VAE (Kingma & Welling, 2022) and incorporating Convolutional Neural Networks (CNNs) as the architecture, and 3DSBDD (Luo et al., 2022) predicts whether the position on grids is occupied with which type of atoms in auto-regressive ways with Graph Neural Networks (GNNs). Then, the development of Equivariant Graph Neural Networks (EGNNs) boosts the SBDD methods to directly generate the continuous 3D positions, such as Pocket2Mol (Peng et al., 2022) generating the molecules’ atom types, positions, and the connected bonds and GraphBP (Liu et al., 2022) employing normalizing flows to generate these attributes, both in an auto-regressive way. The diffusion denoising probabilistic models (DDPM) (Ho et al., 2020) further propel the AI-aided SBDD methods, such as TargetDiff (Guan et al., 2023b), DiffBP (Lin et al., 2022) and DiffSBDD (Schneuing et al., 2022), inspired by EDM (Hoogeboom et al., 2022), generating full atoms’ positions and element types, with different diffusion models. In recent years, domain knowledge has been used to constrain or guide the generation of binding drugs. For example, FLAG (Zhang et al., 2023b) and D3FG (Lin et al., 2023) use prior knowledge of fragment motifs to model the coarse structures, which generate fragments in molecules in one-shot and auto-regressive ways, respectively; And DecompDiff (Guan et al., 2023c) harnesses virtual points searching for arm- and scaffold-clustering as prior knowledge, and using multivariate Gaussian process to model atom positions in different clusters, with validity guidance for sampling. More recently, thanks to technological breakthroughs brought by generative AI, more related approaches have been proposed, which not only build upon previous modeling concepts but also incorporate new generative model techniques. MolCraft (Qu et al., 2024), as the SBDD version of GeoBFN (Song et al., 2024), employs Bayesian Flow Network (Graves et al., 2024) as the variant of diffusion models, to address the continuous-discrete gap in modeling the elements’ type and atoms’ positions by applying continuous noise and smooth transformation; VoxBind (Pinheiro et al., 2024a), following their pioneering exploration of VoxMol (Pinheiro et al., 2024b), continues modeling the 3D voxelization of molecules in a diffusion denosing way, with walk-jump sampling method used to generate molecules.
Table 1: Categorization of included methods. LiGAN, 3DSBDD and VoxBind model atom positions as discrete variables; One-shot methods maintain a constant atom number in a generation, like diffusion-based ones; FLAG and D3FG use fragment motifs, and DecompDiff uses arm&scaffolding priors.
@@
In this way, we can categorize these methods with the three standards:
  • Whether the generation process is auto-regressive or one-shot.
To better articulate, we classify them according to Table. 1, and Figure. 2 gives a simple demonstration of the criteria for taxonomy. A detailed review of these methods is given in Appendix. A.
@@ 3 Task and Dataset
For the de novo molecule generation, we follow the previous protocol to use Crossdocked2020 (Francoeur et al., 2020) and data preparation with splits proposed in LiGAN (Masuda et al., 2020) and 3DSBDD (Luo et al., 2022) as the training and test sets. Besides the de novo generation, the modularized methods can be extended to four subtasks which are branches of our defined fill-in-the-blank of the complex binding graph, including the target-aware linker, fragment, side chain, and scaffold design (Zhang et al., 2024), leading to the generative targets $\mathcal{G}$ for the probabilistic model $p(\mathcal{G}|\mathcal{C},\mathcal{P})$ to be the four components, and the molecular context $\mathcal{C}$ as the rest, as shown in Figure. 3. We demonstrate the significance of the four tasks in lead optimization and how the corresponding datasets are established below. The datasets are all established based on previous splits of Crossdocked2020 for a fair comparison.
Linker Design is a critical strategy in fragment-based drug discovery (Grenier et al., 2023). It focuses on creating linkers that connect two lead fragments and obtaining a complete molecule with enhanced affinity. Effective linker design can significantly influence the efficacy and pharmacokinetics of the drug, by ensuring that the molecule maintains the desired orientation and conformation when bound to its target (Erlanson et al., 2016). We define linkers following specific rules: (i) A linker should act as the intermediate structure connecting two fragments. (ii) A linker should contain at least two atoms on the shortest path between two fragments. (iii) Each connecting fragment must consist of more than five atoms.
Fragment Growing focus on expanding a fragment on the lead compound to better fill the binding pocket (Bancet et al., 2020). It also relates to adjusting pharmacological properties, such as enhancing solubility or reducing toxicity (Hung et al., 2009). We decompose the entire ligand into two fragments and select the smaller one for expansion based on specific criteria: (i) Each fragment must contain more than five atoms. (ii) The smaller fragment should be larger than half the size of the larger one.
Side Chain Decoration differs from fragment growing in that it permits modifications at multiple sites on the lead compound, whereas fragment growing typically modifies a single site. The taxonomy of arms-scaffold in previous work (Guan et al., 2024) is similar to side-chain-scaffold, while we focus on a chemist-intuitive approach, Bemis-Murko decomposition (Bemis & Murcko, 1996), which treats all terminal non-cyclic structures as side chains.
Scaffold Hopping is introduced by  Schneider et al. (1999), as a strategy in medicinal chemistry aiming to replace the core structure of molecules to explore more diverse chemical space or improve specific properties. While various empirical definitions exist (Sun et al., 2012), for consistency, the Bemis-Murko decomposition is utilized to define the scaffold structure.